OpenStudy (anonymous):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

OpenStudy (anonymous):

y sub 2 - y sub 1 / x sub 2 - x sub 1 = -11-13/-4-2 = 4 y= mx+b y= 4x+b since they are just looking for the slope of a line, the answer is simply 4.

OpenStudy (anonymous):

ok thank ya

OpenStudy (anonymous):

how abiout this one

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is the slope of the line passing through the point (4, −2) and perpendicular to the line y = 4x + 2? Write your answer as a fraction.

OpenStudy (anonymous):

oh this one you don't even have to solve

OpenStudy (anonymous):

okay here

OpenStudy (anonymous):

parallel : same slope perpendicular: opposite reciprocal so the answer would be -1/4

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

oh yeah!!!!!!!